# Accelerating Vision Diffusion Transformers with Skip Branches

### v2

@@ 4.2 Main Results
Skip-Cache
@@
In Table 3, we present a quantitative evaluation of all text-to-video models and caching methods. Videos are generated using the prompts from VBench [10], which is considered a more generalized benchmark [47, 10, 48]. Compared with the original Latte, Skip-Cache achieves a comparable VBench score (75.60 vs. 76.14) with only six days of continual pre-training on 330k training samples.
To demonstrate the superiority of the caching mechanism in Skip-Cache, we evaluate two caching settings: caching at timesteps 700–50 and 800–50 (out of 1000 timesteps in total). In both settings, Skip-Cache achieves the highest speedup while maintaining superior scores in PSNR, LPIPS, and SSIM, with only a minor loss in VBench score.
In the first setting, our caching mechanism achieves 1.49$\times$ and 1.72$\times$ speedup with only a 0.12% and 0.22% loss in VBench score, respectively. Among other caching methods, only PAB469 achieves a speedup of more than 1.30$\times$, but at the cost of a 4.50% drop in VBench score. Moreover, our caching method can achieve a 1.90$\times$ speedup while still maintaining absolutely better PSNR, LPIPS, and SSIM scores compared to other caching methods.
Furthermore, in the second setting, we achieve a 2.20$\times$ speedup with just a 1.17% sacrifice in VBench score, representing the highest speedup among current training-free DiT acceleration works.
4.3 Generalize Skip-Cache to image Generation
Hunyuan-DiT [16] is a powerful text-to-image DiT model featuring skip branches, whose effectiveness has been demonstrated in [16]. However, its skip branches have not been explored for accelerating image generation. We leverage these skip branches using the same caching mechanism as Skip-Cache  and compare our caching strategy with other training-free acceleration methods. Furthermore, we extend our proposed Skip-Cache  to class-to-image task in Appendix 6, where Skip-DiT  exceeds vanilla DiT model with only around 38% of its training cost.
@@ 4.3 Generalize Skip-Cache to image Generation
To evaluate the generalization of the caching mechanism in Skip-Cache for text-to-image tasks, we use the zero-shot Fréchet Inception Distance (FID) on the MS COCO [17] 256$\times$256 validation dataset by generating 30,000 images based on its prompts, following the evaluation guidelines established by Hunyuan-DiT. Additionally, we employ Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [43], and Structural Similarity Index Measure (SSIM) [37] to assess the changes introduced by the caching methods. To ensure a fair comparison, we disable the prompt enhancement feature of Hunyuan-DiT. All videos are generated at a resolution of 1024$\times$1024 and subsequently resized to 256$\times$256 for evaluation.
@@
Table 4 provides a comprehensive comparison of Hunyuan-DiT and various caching methods. Notably, our caching mechanism achieves a 2.09$\times$ speedup without any degradation in FID or CLIP scores. Furthermore, it outperforms all other caching methods in terms of PSNR, LPIPS, and SSIM scores, consistently maintaining the highest performance even with a 1.96$\times$ speedup. These findings underscore the robustness and adaptability of our caching mechanism to image generation tasks.
Table 5: Ablation study on different timestep cached. Total timesteps is 1000. During early timesteps (1000$\rightarrow$700), the feature changes rapidly. In the latter timesteps (700$\rightarrow$50) feature dynamics become considerably smoother. We only perform Skip-Cache  in caching timesteps We highlight the best metrics in bold.
@@
Table 6: Compatibility of Skip-DiT with other caching methods. All the methods are performed on Skip-DiT . $n=i$ indicates caching the high-level feature in $x_{t}$ for reuse during the inference of $x_{t-1},x_{t-2},\dots,x_{t-n+1}$. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.
Method VBench (%) $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Skip-DiT 75.60
T-GATE 75.16 24.09 0.16 0.78
$\Delta$-DiT 75.48 22.64 0.17 0.79
FORA 75.38 23.26 0.16 0.79
PAB235 73.79 19.92 0.29 0.68
PAB347 72.08 18.98 0.34 0.65
PAB469 71.64 18.02 0.37 0.62
4.4 Ablation studies
Figure 4: The feature dynamics of Latte with and without Skip Branches. We use cosine similarity to evaluate the differences in features at the same layers across timesteps.
Select the best timesteps to cache
A heat map visualizing feature dynamics across blocks is shown in Figure 4. After incorporating skip branches into Latte, we observe that major changes are concentrated in the early timesteps, with feature dynamics becoming considerably smoother in the later timesteps (700$\rightarrow$50). In contrast, features in Latte exhibit rapid changes across all timesteps. This finding highlights that caching during smoother timesteps leads to significantly better performance, supporting the hypothesis that smooth features enhance caching efficiency in DiT. In Table 5, we further validate this observation: under equivalent throughput, caching in the later timesteps (700$\rightarrow$50), where features are smoother, outperforms caching in the earlier timesteps (950$\rightarrow$300), achieving superior PSNR, LPIPS, and SSIM scores. Additionally, we segmented the rapidly changing timesteps and experimented with three caching ranges: 900$\rightarrow$50, 800$\rightarrow$50, and 700$\rightarrow$50. The results show that increasing the ratio of smoother regions significantly improves caching performance. These findings underscore the importance of leveraging smoother feature dynamics for optimal caching.
Compatibility of Skip-DiT  with other caching methods
As shown in Table 6, we extend the existing DiT caching methods to Skip-DiT and observe slight performance improvements. Specifically, in $\Delta$-DiT, the middle blocks are cached instead of the front blocks, as discussed in Section 4.1, to leverage Skip-DiT’ U-shaped structure. Taking PAB [47] as an example, it loses 1.15% less in VBench score and achieves noticeably better PSNR and SSIM scores on Skip-DiT compared to Latte, highlighting the potential of Skip-DiT to enhance cache-based methods, and proving the superior caching efficiency of the model with better feature smoothness.
@@ 5 Conclusion
In this work, we introduce Skip-DiT, a skip-branch-enhanced DiT model designed to produce smoother features and propose Skip-DiT cache to improve caching efficiency in video and image generation tasks. By enhancing feature smoothness across timesteps, Skip-DiT unlocks the potential to cache most blocks while maintaining high-generation quality. Additionally, Skip-Cache leverages its U-net-style architecture to enable cross-timestep feature caching. Our approach achieves maximum speedup in cache-based visual DiT generation while preserving the highest similarity to original outputs. Furthermore, we analyze feature dynamics before and after incorporating skip branches, demonstrating the effectiveness of caching at timesteps with smoother features. We also show that Skip-DiT is compatible with other caching methods, further extending its applicability. Overall, Skip-DiT can seamlessly integrate with various DiT backbones, enabling real-time, high-quality video and image generation while consistently outperforming baseline methods. We believe Skip-Cache offers a simple yet powerful foundation for advancing future research and practical applications in visual generation.
@@ 6 Class-to-image Generation Experiments
Peebles and Xie [22] proposed the first diffusion model based on the transformer architecture, and it outperforms all prior diffusion models on the class conditional ImageNet [7] 512$\times$512 and 256$\times$256 benchmarks. We add skip branches to its largest model DiT-XL/2 to get Skip-DiT. We train Skip-DiT on class conditional ImageNet with resolution 256$\times$256 from scratch with completely the same experiments setting as DiT-XL/2, and far exceeds DiT-XL/2 with only around 38% of its training cost.
@@
We modify the structure of DiT-XL/2 following the methodology outlined in Section 3 and train Skip-DiT  for 2,900,000 steps on 8 A100 GPUs, compared to 7,000,000 steps for DiT-XL/2, which also uses 8 A100 GPUs. The datasets and other training settings remain identical to those used for DiT-XL/2, and we utilize the official training code of DiT-XL/2***https://github.com/facebookresearch/DiT. The performance comparison is presented in Table 7, which demonstrates that Skip-DiT  significantly outperforms DiT-XL/2 while requiring only 38% of its training steps, highlighting the training efficiency and effectiveness of Skip-DiT.
Table 7: Comparison of training efficiency between DiT-XL/2 and Skip-DiT . Images are generated with a 250-step DDPM solver. The term cfg refers to classifier-free guidance scales, where metrics for cfg=1.0 are computed without classifier-free guidance. The best metrics are highlighted in bold. Skip-DiT  significantly exceeds DiT-XL/2 with much less training steps.
@@
Skip-Cache
@@
Skip-Cache
@@
We evaluate Skip-Cache on Skip-DiT  and compare its performance against two other caching methods: $\Delta$-DiT and FORA. As shown in Table 8, Skip-Cache achieves a 1.46$\times$ speedup with only a minimal FID loss of $0.02$ when the classifier-free guidance scale is set to 1.5, compared to the 7–8$\times$ larger losses observed with $\Delta$-DiT and FORA. Moreover, even with a 1.9$\times$ acceleration, Skip-DiT  performs better than the other caching methods. These findings further confirm the effectiveness of Skip-DiT  for class-to-image tasks.

### v4

@@ 4.3 Results on Image Generation Tasks
Skip-DiT  with Cached Inference
@@
Table 4: Class-to-image generation performance. Speedups are calculated on an H100 GPU with a sample batch size of 8. The definition of $n=i$ is the same as other tables. The classifier-free guidance scale is set to 1.5. Skip-DiTis trained with 2.9M steps. We highlight baseline models in grey and the best metrics in bold.
DiT-XL
original 2.30 4.71 276.26 82.68 57.65 1.00$\times$
FORA 2.45 5.44 265.94 81.21 58.36 1.57$\times$
Delta-DiT 2.47 5.61 265.33 81.05 58.83 1.45$\times$
Skip-DiT  with Cached Inference
original 2.29 4.58 281.81 82.88 57.53 1.00$\times$
$n=2$ 2.31 4.76 277.51 82.52 58.06 1.46$\times$
$n=3$ 2.40 4.98 272.05 82.14 57.86 1.73$\times$
$n=4$ 2.54 5.31 267.34 81.60 58.31 1.93$\times$
We extend Skip-DiT to the class-to-image task, where Skip-DiT exceeds DiT-XL with only around 23% of its training cost (1.6M training steps vs.7M), and we choose the Skip-DiT trained with 2.9M steps for evaluation. Hunyuan-DiT [14] is a powerful text-to-image DiT model featuring LSCs. However, its skip connections have not been fully explored to accelerate generation. We first explore their impact and efficiency potential and integrate it with our caching mechanism without architectural modifications.
@@
To evaluate the generalization of the caching mechanism in Skip-DiT for text-to-image tasks, we use the zero-shot Fréchet Inception Distance (FID) on the MS-COCO [15] validation dataset by generating 30k images based on its prompts, following the guidelines of Hunyuan-DiT. Additionally, we also employ Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS), and Structural Similarity Index Measure (SSIM) to evaluate similarity. For the class-to-image task, we evaluate models on ImageNet [7] by generating 50k images, and calculating evaluation metrics used in Peebles and Xie [22].
@@
Table 3 comprehensively compares Hunyuan-DiT and various caching methods. Notably, our caching mechanism achieves a 2.09$\times$ speedup without any degradation in FID or CLIP scores. Furthermore, it outperforms all other caching methods in terms of PSNR, LPIPS, and SSIM scores, consistently maintaining the highest performance even with a 1.96$\times$ speedup. These findings underscore the robustness and adaptability of our caching mechanism to image generation tasks. Additionally, the unusual improvement in FID and CLIP scores is due to: ❶ In HunyuanDiT’s standard evaluation pipeline, FID fluctuates due to zero-shot evaluation and resolution scaling ($1024\rightarrow 256$) prior to evaluation. ❷ The CLIP score is fluctuating within a reasonable range (-0.16, 0.2). These fluctuations are also observed in $\Delta$-DiT. The results on the class-to-image task shown in Table 4, Skip-DiT  also achieve the least loss while maintaining higher speedup.
4.4 Ablation studies
Identify Dynamic Phase for Better Caching
After incorporating LSCs into DiT, we can statically choose the timesteps for caching with feature stability, thus we can identify the dynamic phase by caching similarity shown in Figure 3. However, previous work has observed that the denoising process is split into different phases in U-Net diffusion models by Zhang et al. [49] and  Li et al. [13], highlighting the hidden dynamic phases which may lead to better performance. A heat map visualizing inner feature dynamics within blocks is shown in Figure 5, and significant changes are observed in the early denoising phase. In Table 5, we further validate this observation: under equivalent throughput, caching in the later timesteps (700$\rightarrow$50), outperforms caching in the earlier timesteps (950$\rightarrow$300), achieving superior PSNR, LPIPS, and SSIM scores. Additionally, we segmented the rapidly changing timesteps and experimented with additional caching ranges. All of these results show that increasing the ratio of the former phase significantly improves caching performance.
Table 5: Ablation study on different timestep ranges for caching. Total timesteps is 1000. We only perform cache in Skip-DiT  at caching timesteps. We highlight the best metrics in bold.
@@ 4.4 Ablation studies
Figure 5: The feature dynamics of Latte with LSCs. Differences in features at the same layers across timesteps are evaluated.
Compatibility of Skip-DiT
As shown in Table 6, we extend the other text-to-video DiT caching methods to Skip-DiT and observe slight performance improvements. Taking PAB [53] as an example, it loses 1.15% less in VBench and achieves noticeably better PSNR and SSIM scores on Skip-DiT , demonstrating the compatibility of Skip-DiT , and that the stable features in Skip-DiT enhance the model robustness and caching efficiency.
Table 6: Compatibility of Skip-DiT  with other caching methods. The performance or other caching methods on vanilla DiT are shown in the table 2. Obvious improvements can be seen especially in PAB. We highlight the best metrics in bold.
Method $\Delta$VBench (%) $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
T-GATE 0.44 24.09 0.16 0.78
$\Delta$-DiT 0.12 22.64 0.17 0.79
FORA 0.22 23.26 0.16 0.79
PAB235 1.81 19.92 0.29 0.68
PAB347 3.52 18.98 0.34 0.65
PAB469 3.96 18.02 0.37 0.62
Analysis of Block Selection for Caching
To identify the optimal block for caching, we analyze the feature similarity across timesteps in the final three blocks of Latte-T2V (Table 7). Our analysis reveals that caching at block 27, the last DiT block, yields the best performance. This block, which is connected to block 0 via the primary Long Skip Connection (LSC), not only exhibits maximum feature similarity but also enables the highest speedup. These results demonstrate the superior caching efficiency achieved by Skip-DiT.
Table 7: Connection selection for caching in the text-to-video task. The best metrics are emphasized in bold.
LSC Cached Similarity(%$\uparrow$ VBench(%) $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
0 [1,26] 99.88 75.51 29.52 0.06 0.89
1 [2,25] 99.86 75.44 29.65 0.06 0.89
2 [3,24] 99.66 75.49 29.44 0.06 0.88
Comparison in Few-Step Generation
We benchmarked Skip-DiT against a vanilla DiT in a few-step text-to-video generation scenario, using a 5 and 10-step DDIM scheduler instead of the standard 50 steps. As shown in Table 8, while the vanilla DiT holds a slight advantage in the SSIM metric, Skip-DiT achieves superior visual quality and semantic consistency. These results again validate the balanced and robust performance of our architecture, even in a highly accelerated, few-step setting.
Table 8: Text-to-video generation performance under the few-step scenario. The best metrics are emphasized in bold.
NFE Model $\Delta$VBench(%) PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
10 steps Latte $\downarrow$1.73 15.96 0.47 0.61
Skip-DiT $\downarrow$0.94 16.18 0.43 0.59
5 steps Latte $\downarrow$9.40 12.94 0.73 0.53
Skip-DiT $\downarrow$8.55 13.17 0.69 0.50
@@ 5 Conclusion
In this work, we identify the Dynamic Feature Instability in vanilla DiT models and introduce Skip-DiT , a Long-Skip-Connections enhanced DiT model to produce more stable features and build training and inference efficient diffusion transformers. We designed a simpler but more efficient caching mechanism for Skip-DiT  to improve inference efficiency in video and image generation tasks. By improving the stability of dynamic features, Skip-DiT  unlocks the potential to cache most blocks while maintaining high generation quality. In addition, we guarantee stability and efficiency with theoretical analysis from the perspective of the spectral norm. Our proposed DiT achieves significant speedup in training convergence and model capability. With the caching mechanism, we achieve maximum speedup in cache-based acceleration while maintaining the highest similarity to the original outputs. Furthermore, we identify the dynamic phase of Skip-DiT  for a better caching strategy, demonstrating the effectiveness of caching. We also show that Skip-DiT  is compatible with other caching methods, further extending its applicability. Overall, Skip-DiT  can be seamlessly integrated with various DiT backbones, enabling real-time, high-quality video and image generation, while consistently outperforming baseline methods. We believe that Skip-DiT  provides a simple yet powerful foundation for advancing future research and applications.
6 Limitations
Despite achieving significant speedups, Skip-DiT  inherits limitations from its DiT foundation. including high computational demands, reliance on large-scale training data, and quadratic complexity that hinders high-resolution generation. Furthermore, Skip-DiT  introduces a marginal parameter overhead (5.5% for class-to-image, 3.5% for text-to-video) and extra GPU memory (scales with the batch size $\mathcal{N}$), requiring an additional $\mathcal{N}\times 0.37\%$ for the class-to-image model and $\mathcal{N}\times 4.55\%$ for the text-to-video model. These models remain feasible for deployment with proper $\mathcal{N}$.
Acknowledgement
During this project, Guanjie Chen is supported by Shanghai Artificial Intelligence Laboratory.
@@ Acknowledgement
7 Detailed Proof of Theorems
Consider an ideal denoising diffusion transformer $M$ with $L$ identical blocks, where the denoising capability induces the following fundamental properties:
  • Noise Reduction Invariance: Each transformer block $\mathcal{T}_{l}$ strictly reduces noise magnitude of input $h$. This directly implies the Jacobian spectral norm ($\sigma_{\max}$) constraint:
    $\sigma_{\max}\left(\frac{\partial\mathcal{T}_{l}}{\partial h}\right)\triangleq\gamma_{l}<1$ (9)
  • Transformer Blocks Homogeneity: Identical noise reduction ratio across layers $\gamma_{l}=\gamma,\quad\forall l\in\{1,...,L\}$. Thus, the complete model satisfies
    $\sigma_{\max}(M)=\prod_{l=1}^{L}\gamma=\gamma^{L}\ll 1$ (10)
7.1 Theorem1
The spectral norm of the Jacobian matrix of DiT with Long-Skip-Connections is controlled tighter than that of Vanilla DiT $M$, making the Skip-DiT model more robust, numerically stable, and capable of converging faster.
Proof.
Define layer transformations for $L/2<l\leq L$ and their Jacobian matrices, where $\mathcal{T}$ denotes the Transformer block and $10<\alpha<$ 1:
$h_{l+1}^{\text{vanilla}}=\mathcal{T}(h_{l}^{\text{vanilla}}),h_{l+1}^{\text{skip}}=(1-\alpha)\cdot\mathcal{T}(h_{l}^{\text{skip}})+\alpha\cdot h_{L-l}^{\text{skip}}.$
The corresponding Jacobian matrices $J$ are:
$J_{l}^{\text{vanilla}}=\frac{\partial\mathcal{T}\left(h_{l}\right)}{\partial h_{l}},J_{l}^{\text{skip}}=(1-\alpha)\cdot\frac{\partial\mathcal{T}\left(h_{l}\right)}{\partial h_{l}}+\alpha\cdot\frac{\partial h_{L-l}}{\partial h_{l}}.$
Applying subadditivity and submultiplicativity of spectral norms $\sigma_{max}$:
$\sigma_{\max}(J_{l}^{\mathrm{skip}})$ $\leq(1-\alpha)\gamma+\alpha\gamma^{2l-L}.$ (11)
Given $\gamma<1$ and $2l-L\geq 1$ for $l>L/2$, we establish the layer-wise bound:
$\sigma_{\max}(J_{l}^{\mathrm{skip}})<\gamma=\sigma_{\max}(J_{l}^{\mathrm{vanilla}}).$ (12)
For the full model Jacobian $J=\prod_{l=1}^{L}J_{l}$, the spectral norms satisfy:
$\sigma_{\max}(J^{\mathrm{skip}})$ $\leq\prod_{l=0}^{L-1}\left[(1-\alpha)\gamma+\alpha\gamma^{2l-L}\right]$
$<\prod_{l=0}^{L-1}\gamma=\gamma^{L}=\sigma_{\max}(J^{\mathrm{vanilla}}).$ (13)
This spectral radius reduction attenuates perturbation growth through $\|\delta h_{L}\|\leq(\sigma_{\max}(J))^{L}\|\delta h_{0}\|$, stabilizes gradient flow via $\|\nabla_{\theta}\mathcal{L}\|\leq\sigma_{\max}(J)\|\nabla_{h_{L}}\mathcal{L}\|$, and improves the Lipschitz constant $\mathrm{Lip}(M_{\mathrm{skip}})<\mathrm{Lip}(M_{\mathrm{vanilla}})$. ∎
7.2 Theorem2
Under feature reuse with interval $\tau$, Skip-DiT achieves tighter perturbation bounds than Vanilla DiT, enabling larger allowable reuse intervals while maintaining error tolerance $\epsilon_{\mathrm{max}}$
Proof.
Let $h_{t}$ be the reused feature with error $\epsilon_{t}$, $\delta$ the perturbation bound per step, $x$ the input, and $\theta$ the model parameters. Through the differential mean-value theorem and error propagation:
$\|f(x;\theta+\Delta\theta)-f(x;\theta)\|$ $\leq\sup_{\theta^{\prime}}\left\|\frac{\partial f}{\partial\theta}\right\|\|\Delta\theta\|\triangleq\delta$ (14)
$\epsilon_{t+1}=\|h_{t+1}-h_{t}\|$ $\leq\mathrm{Lip}(M)\epsilon_{t}+\delta$ (15)
From Theorem 1, the Lipschitz constants satisfy:
$\mathrm{Lip}(M_{\mathrm{skip}})=\sigma_{\max}(J^{\mathrm{skip}})<\gamma^{L}=\mathrm{Lip}(M_{\mathrm{vanilla}})$ (16)
The perturbation bounds inherit:
$\delta_{\mathrm{skip}}=C_{\mathrm{skip}}\|\Delta\theta\|<C_{\mathrm{vanilla}}\|\Delta\theta\|=\delta_{\mathrm{vanilla}}$ (17)
where $C_{\cdot}$ bounds the parameter-to-output Jacobian.
$\left\|\frac{\partial f_{\mathrm{skip}}}{\partial\theta}\right\|$ $=\left\|\prod_{l=1}^{L}\frac{\partial h_{l}}{\partial\theta}\right\|\leq\prod_{l=1}^{L}\left\|\frac{\partial h_{l}}{\partial\theta}\right\|$
$=\prod_{l=1}^{L}\left[(1-\alpha)\gamma+\alpha\gamma^{2l-L}\right]\cdot C_{\mathrm{base}}$
$<\prod_{l=1}^{L}\gamma\cdot C_{\mathrm{base}}=C_{\mathrm{vanilla}}$
For $T$-step feature reuse, cumulative error develops as:
$\epsilon_{T}\leq\frac{\mathrm{Lip}(M)^{T}-1}{\mathrm{Lip}(M)-1}\delta$ (18)
Given $\sigma_{\max}(J^{\mathrm{skip}})<\gamma^{L}$ and $\delta_{\mathrm{skip}}<\delta_{\mathrm{vanilla}}$, the maximal interval $\tau$ satisfies:
$\frac{\sigma_{\max}(J^{\mathrm{skip}})^{\tau}-1}{\sigma_{\max}(J^{\mathrm{skip}})-1}\delta_{\mathrm{skip}}=\frac{\gamma^{L\tau}-1}{\gamma^{L}-1}\delta_{\mathrm{vanilla}}=\epsilon_{\mathrm{max}}$ (19)
Solving equation 19 yields $\tau_{\mathrm{skip}}>\tau_{\mathrm{vanilla}}$ under identical $\epsilon_{\mathrm{max}}$, proving Skip-DiT permits larger reuse intervals. ∎
@@ 8 Class-to-image Generation Experiments
Peebles and Xie [22] proposed the first diffusion model based on the transformer architecture, and it outperforms all prior diffusion models on the class conditional ImageNet [7] 512$\times$512 and 256$\times$256 benchmarks. We add skip connections to its largest model, DiT-XL, to get Skip-DiT. We train Skip-DiT on class conditional ImageNet with resolution 256$\times$256 from scratch with completely the same experimental settings as DiT-XL, and far exceeds DiT-XL with only around 38% of its training cost.
@@
We add long-skip-connections in DiT-XL and train Skip-DiT  for 2.9M steps on 8 A100 GPUs, compared to 7M steps for DiT-XL, which also uses 8 A100 GPUs. The datasets and other training settings remain identical to those used for DiT-XL, and we utilize the official training code of DiT-XLhttps://github.com/facebookresearch/DiT. The performance comparison is presented in Table 9, which demonstrates that Skip-DiT  significantly outperforms DiT-XL while requiring only 23% of its training steps (1.6M vs. 7M), highlighting the training efficiency and effectiveness of Skip-DiT.
Table 9: Comparison of training efficiency between DiT-XL and Skip-DiT . Images are generated with a 250-step DDPM solver. The term cfg refers to classifier-free guidance scales, where metrics for cfg=1.0 are computed without classifier-free guidance. The best metrics are highlighted in bold. Skip-DiT  significantly exceeds DiT-XL/2 with much less training steps.
@@
Faster-Diff 4.96 10.19 223.21 75.21 59.28 1.42$\times$
Skip-DiT with Cached Inference
@@
Faster-Diff 22.98 18.09 80.41 56.53 67.41 1.42$\times$
Skip-DiT with Cached Inference
@@
We evaluate Skip-DiT  and compare its performance against two other caching methods: $\Delta$-DiT and FORA. As shown in Table 10, Skip-DiT achieves a 1.46$\times$ speedup with only a minimal FID loss of $0.02$ when the classifier-free guidance scale is set to 1.5, compared to the 7–8$\times$ larger losses observed with $\Delta$-DiT and FORA. Moreover, even with a 1.9$\times$ acceleration, Skip-DiT  performs better than the other caching methods. These findings further confirm the effectiveness of Skip-DiT  for class-to-image tasks.